An electronic starts from rest and falls through a potential rise of 80 v. What is its final speed?
1
Expert's answer
2019-01-10T15:36:26-0500

We can find the final speed of the electron from the work-kinetic energy theorem. It states that the work done by potential difference is equal to the change in kinetic energy of the electron:

"\u2206PE=KE_f - KE_i,""qV = \frac{1}{2}mv_f^2 - 0,"

here,

"q"

is the charge of the electron,

"V"

is the potential difference,

"m"

is the mass of the electron and

"v_f"

is the final speed of the electron.

Then, from this formula we can find the final speed of the electron:

"v_f = \sqrt{\frac{2qV}{m}}."

Let's substitute the numbers:

"v_f = \sqrt{\frac{2 \cdot 1.6 \cdot 10^{-19} C \cdot 80 V}{9.11 \cdot 10^{-31} kg}} =5.3 \cdot 10^6 \frac{m}{s}."

Answer:

"v_f = 5.3 \cdot 10^6 \frac{m}{s}"

.